OpenStudy (anonymous):

GIVING MEDALS if a=22, b=14, and c=30, find the area of triangle ABC

OpenStudy (kc_kennylau):

HERON'S FORMULA OP :D s=0.5(22+14+30)=33

OpenStudy (kc_kennylau):

Area \[=\sqrt{s(s-a)(s-b)s-c)}\] \[=\sqrt{33*11*19*3}\]

OpenStudy (kc_kennylau):

\[=33\sqrt{19}\]

OpenStudy (anonymous):

is that the answer?!

OpenStudy (kc_kennylau):

yep

OpenStudy (kc_kennylau):

just checked with wolfram alpha, which agrees with me

OpenStudy (anonymous):

You can use Heron's formula, which is: sq rt [s(s-a)(s-b)(s-a)], where s = (a+b+c)/2 From there, you plug in your numbers. s = (22+14+30)/2 s = 33 Then you plug s into Heron's formula. = sq rt [s(s-a)(s-b)(s-a)] = sq rt [33(33-22)(33-14)(33-30)] = sq rt [33(11) (19) (3)] = sq rt [20691] = 143.84 The answer is D.

OpenStudy (kc_kennylau):

where is the choices given?

OpenStudy (kc_kennylau):

are*

OpenStudy (anonymous):

its not multiple choice

OpenStudy (kc_kennylau):

well i said the answer is 33sqrt(19), which is approximately equivalent to 143.84